# Neural network learns low-dimensional polynomials with SGD near the information-theoretic limit

### v1

@@ B.2 Random Initialization
In Section B.3.1 we focus on the neurons with slightly larger initial alignment that satisfy $\kappa^{0}_{j}=\boldsymbol{\theta}^{\top}\boldsymbol{w}^{0}_{j}\geq 2C_{2}d^{-\frac{1}{2}}$ at initialization, where constant $C_{2}$ grows at most polylogarithmically in $d$. The following lemma states that roughly a constant portion of the neurons satisfies this initial alignment condition.
@@
Proof of Lemma 7. Because $\kappa^{0}=v^{\top}w\overset{\mathrm{d}}{=}\frac{{\sf e}_{1}^{\top}g}{g}$, where $g\sim\mathcal{N}(0,\boldsymbol{I}_{d})$,
$\mathbb{P}[\kappa^{0}_{j}\geq 2C_{2}d^{-\frac{1}{2}}]$ $=\mathbb{P}_{g\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}{\sf e}_{1}^{\top}g\geq 4C_{2}\land\|g\|\leq 2d^{\frac{1}{2}}\bigg{]}$ (B.10)
$\geq\mathbb{P}_{g\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}{\sf e}_{1}^{\top}g\geq 4C_{2}\bigg{]}-\mathbb{P}_{g\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\|g\|\geq 2d^{\frac{1}{2}}\bigg{]}$ (B.11)
$\gtrsim\frac{\sqrt{2e(\beta-1)}}{2\beta\sqrt{\pi}}e^{-8\beta C_{2}^{2}}-e^{-\Omega(d)}.$ (B.12)
@@
By letting $\beta=2$, we have that $\mathbb{P}[\kappa^{0}_{j}\geq C_{2}d^{-\frac{1}{2}}]\gtrsim e^{-16C_{2}^{2}}.$ Due to symmetry, $\mathbb{P}[\kappa^{0}_{j}\leq 2C_{2}d^{-\frac{1}{2}}]=\mathbb{P}[\kappa^{0}_{j}\geq 2C_{2}d^{-\frac{1}{2}}]$. ∎
B.3 Population Update
@@
We first analyze the training of first-layer parameters by evaluating the expected (population) update of two gradient steps with the same training example. At each step, the parameters are updated as
$=\boldsymbol{w}^{t}_{j}-\eta^{t}\tilde{\nabla}_{\boldsymbol{w}}\Bigg{(}\frac{1}{N}\sum_{j=1}^{N}a_{j}\sigma({\boldsymbol{w}_{j}^{t}}^{\top}\boldsymbol{x})\Bigg{)}^{2}+2\eta^{t}\tilde{\nabla}_{\boldsymbol{w}}\Bigg{(}y\frac{1}{N}\sum_{j=1}^{N}a_{j}\sigma({\boldsymbol{w}_{j}^{t}}^{\top}\boldsymbol{x})\Bigg{)}.$ (B.14)
While the second term scales with $\eta^{t}a_{j}^{2}=\eta^{t}c_{a}^{2}$, the third term scales with $\eta^{t}a_{j}=\eta^{t}c_{a}$. Thus, by setting the second-layer scale $c_{a}$ sufficiently small, we can ignore the interaction of neurons; similar mechanism also appeared in [BES+22, AAM23]. Specifically, in the following, we show that the strength of the signal is $\boldsymbol{\theta}^{\top}\boldsymbol{w}_{j}^{t}\gtrsim d^{-\frac{1}{2}}$. Thus, by simply letting $c_{a}\lesssim C_{3}^{-1}d^{-\frac{1}{2}}$, we can ignore the effect of the squared term. Thus we may focus on the following correlational update:
$\boldsymbol{w}^{t+1}_{j}\leftarrow\boldsymbol{w}^{t}_{j}+\eta^{t}\tilde{\nabla}_{\boldsymbol{w}}\Bigg{(}y\frac{1}{N}\sum_{j=1}^{N}a_{j}\sigma({\boldsymbol{w}_{j}^{t}}^{\top}\boldsymbol{x})\Bigg{)}.$ (B.15)
Due to the absence of interaction between neurons, we omit the subscript $j$ for the index of neurons and ignore the prefactor of $N$ (which can be absorbed into the learning rate); multiplying $N$ to $\eta^{t}$ specified below recovers the scaling of $\eta^{t}$ presented in the main text.
B.3.1 When $\mathrm{IE}(\sigma_{*}^{I})=2$ with $I\geq 2$
First we consider the most technically difficult case, when $\mathrm{IE}[\sigma_{*}]\geq 3$ and the information exponent cannot be lowered to 1 for $i\leq C_{\sigma}$; in this case, from Proposition 5 we know that there exists some $2\leq i\leq C_{\sigma}$ such that $\mathrm{IE}[\sigma_{*}^{i}]=2$ and we let $I$ be the first such $i$.

### v2

@@ B.2 Initialization
We first consider the initial alignment. In the following sections, we focus on the neurons that satisfy $\kappa^{0}_{j}=\boldsymbol{\theta}^{\top}\boldsymbol{w}^{0}_{j}\geq 2c_{\eta}^{-1}d^{-\frac{1}{2}}$ at the initialization. The following lemma states that roughly a constant portion of the neurons satisfy the initial alignment condition upon random initialization. In particular, if we take $c_{\eta}=\Omega((\log\log d)^{-\frac{1}{2}})$, the fraction of neurons that satisfy the initial alignment condition is at least $e^{-16c_{\eta}^{-2}}=\tilde{\Omega}(1)$. Let us write $C_{2}=c_{\eta}^{-1}$ for simplicity in the following.
@@
Proof of Lemma 14. Because $\kappa^{0}=\boldsymbol{v}^{\top}\boldsymbol{w}\overset{\mathrm{d}}{=}\frac{{\boldsymbol{e}}_{1}^{\top}\boldsymbol{g}}{\|\boldsymbol{g}\|}$, where $\boldsymbol{g}\sim\mathcal{N}(0,\boldsymbol{I}_{d})$,
$\mathbb{P}[\kappa^{0}_{j}\geq 2C_{2}d^{-\frac{1}{2}}]$ $=\mathbb{P}_{\boldsymbol{g}\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}{\boldsymbol{e}}_{1}^{\top}\boldsymbol{g}\geq 4C_{2}\land\|\boldsymbol{g}\|\leq 2d^{\frac{1}{2}}\bigg{]}$ (B.17)
$\geq\mathbb{P}_{\boldsymbol{g}\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}{\boldsymbol{e}}_{1}^{\top}\boldsymbol{g}\geq 4C_{2}\bigg{]}-\mathbb{P}_{\boldsymbol{g}\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\|\boldsymbol{g}\|\geq 2d^{\frac{1}{2}}\bigg{]}$ (B.18)
$\gtrsim\frac{\sqrt{2e(\beta-1)}}{2\beta\sqrt{\pi}}e^{-8\beta C_{2}^{2}}-e^{-\Omega(d)},$ (B.19)
@@
where we used Lemma 15 for the final inequality. By letting $\beta=2$, we have that $\mathbb{P}[\kappa^{0}_{j}\geq C_{2}d^{-\frac{1}{2}}]\gtrsim e^{-16C_{2}^{2}}.$ Because of the symmetry, $\mathbb{P}[\kappa^{0}_{j}\leq 2C_{2}d^{-\frac{1}{2}}]=\mathbb{P}[\kappa^{0}_{j}\geq 2C_{2}d^{-\frac{1}{2}}]$. ∎
B.3 Weak Recovery: Population Update
@@
We divide the first layer training into the first phase (weak recovery) and the second phase (strong recovery). We first evaluate the expected update of two gradient steps with the same training example.
Lemma 16.
Let $\eta^{2t},\eta^{2t+1}=\eta=c_{\eta}d^{-1}$, $\xi^{2(t+1)}=\xi=1-c_{\xi}d^{-\frac{(p_{*}-2)_{+}}{2}}$. Suppose that the link function satisfies $\mathrm{IE}(\sigma_{*}^{I})=\mathrm{GE}(\sigma_{*})=p_{*}$ (we choose the smallest such $I$) and activation functions satisfy all of the assumptions in Section B.1 for weak recovery. Then, for $\boldsymbol{w}^{2t}$ with $c_{\eta}^{-1}d^{-\frac{1}{2}}\leq\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2t}\leq c_{\eta}^{I}$, the alignment $\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2(t+1)}$ can be evaluated as,
$\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2(t+1)}\geq\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2t}+c_{\eta}^{I}c_{\xi}c_{\sigma}d^{-\frac{p_{*}}{2}\lor 1}(\kappa^{2t})^{p_{*}-1}+c_{\eta}c_{\xi}d^{-\frac{p_{*}}{2}\lor 1}\nu^{2t}.$ (B.20)
Here $c_{\sigma}=p_{*}!\alpha_{p_{*}}\beta_{p_{*}}$ (when $\mathrm{IE}(\sigma_{*})=\mathrm{GE}(\sigma_{*})$) or $c_{\sigma}=\frac{p_{*}!H(\sigma_{*}^{I};p_{*})H(\sigma^{(I)}(\sigma^{(1)})^{I-1};p_{*}-1)}{2(I-1)!}$ (otherwise), and $\nu^{2t}$ is a mean-zero sub-exponential random variable.
Proof.  The expected alignment $\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2(t+1)}$ after two gradient steps from $\boldsymbol{w}^{2t}=\boldsymbol{\omega}$ using the same sample $(\boldsymbol{x},y)$, step size $\eta^{2t}=\eta^{2t+1}=\eta=c_{\eta}d^{-1}$ and momentum parameter $\xi^{2(t+1)}=\xi=1-c_{\xi}d^{-\frac{(p_{*}-2)_{+}}{2}}$ is evaluated as follows. With a projection matrix $\boldsymbol{P}_{\boldsymbol{\omega}}=\boldsymbol{I}-\boldsymbol{\omega}\boldsymbol{\omega}^{\top}$, the first step updates the weight as
$\boldsymbol{w}^{2t+1}\leftarrow\boldsymbol{w}^{2t}+\eta\tilde{\nabla}_{\boldsymbol{w}}y\sigma({\boldsymbol{w}^{2t}}^{\top}\boldsymbol{x})=\boldsymbol{\omega}+\eta y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x},$ (B.21)
and the next gradient step with the same sample is computed as
~$\tilde{\nabla}_{\boldsymbol{w}}y\sigma({\boldsymbol{w}^{2t+1}}^{\top}\boldsymbol{x})$ $=y\sigma^{\prime}({\boldsymbol{w}^{2t+1}}^{\top}\boldsymbol{x})\boldsymbol{x}$ (B.22)
$=y\sigma^{\prime}\big{(}(\boldsymbol{\omega}+\eta y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x})^{\top}\boldsymbol{x}\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}$ (B.23)
$=y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x},$ (B.24)
here we used the notation $\|\boldsymbol{\theta}\|_{A}^{2}=\boldsymbol{\theta}^{\top}A\boldsymbol{\theta}$ for a vector $\boldsymbol{\theta}\in\mathbb{R}^{d}$ and a positive symmetric matric $A\in\mathbb{R}^{d\times d}$. From (B.21) and (B.24), the parameter after the two steps is obtained as
$\boldsymbol{w}^{2(t+1)}$ $\leftarrow\boldsymbol{w}^{2t+1}+\eta\tilde{\nabla}_{\boldsymbol{w}}y\sigma({\boldsymbol{w}^{2t+1}}^{\top}\boldsymbol{x})$ (B.25)
$=\boldsymbol{\omega}+\eta y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}+\eta y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}.$ (B.26)
$=\boldsymbol{\omega}+\eta\boldsymbol{g}^{2t},$ (B.27)
where
$\boldsymbol{g}^{2t}=y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}+y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}.$ (B.28)
Finally, the normalization step yields
$\boldsymbol{w}^{2(t+1)}$ $\leftarrow\frac{\boldsymbol{w}^{2(t+1)}-\xi^{2(t+1)}(\boldsymbol{w}^{2(t+1)}-\boldsymbol{w}^{2t})}{\|\boldsymbol{w}^{2(t+1)}-\xi^{2(t+1)}(\boldsymbol{w}^{2(t+1)}-\boldsymbol{w}^{2t})\|}=\frac{\boldsymbol{\omega}+\eta\xi\boldsymbol{g}^{2t}}{\|\boldsymbol{\omega}+\eta\xi\boldsymbol{g}^{2t}\|}=\frac{\boldsymbol{\omega}+c_{\eta}c_{\xi}d^{-\frac{p_{*}}{2}\lor 1}\boldsymbol{g}^{2t}}{\|\boldsymbol{\omega}+c_{\eta}c_{\xi}d^{-\frac{p_{*}}{2}\lor 1}\boldsymbol{g}^{2t}\|}.$ (B.29)
Therefore, by writing $\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2t}=\kappa^{2t}$, the update of the alignment is
$\kappa^{2(t+1)}=\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2(t+1)}$ (B.30)
$=\frac{\kappa^{2t}+c_{\eta}c_{\xi}d^{-\frac{p_{*}}{2}\lor 1}\boldsymbol{\theta}^{\top}\boldsymbol{g}^{2t}}{\|\boldsymbol{\omega}+c_{\eta}c_{\xi}d^{-\frac{p_{*}}{2}\lor 1}\boldsymbol{g}^{2t}\|}$ (B.31)
$\geq\kappa^{2t}+c_{\eta}c_{\xi}d^{-\frac{p_{*}}{2}\lor 1}\boldsymbol{\theta}^{\top}\boldsymbol{g}^{2t}-\frac{1}{2}\kappa^{2t}c_{\eta}^{2}c_{\xi}^{2}d^{-p_{*}\lor 2}\|\boldsymbol{g}^{2t}\|^{2}-\frac{1}{2}c_{\eta}^{3}c_{\xi}^{3}d^{-\frac{3p_{*}}{2}\lor 3}|\boldsymbol{\theta}^{\top}\boldsymbol{g}^{2t}|\|\boldsymbol{g}^{2t}\|^{2}.$ (B.32)
We can easily see that $\mathbb{E}[\|\boldsymbol{g}^{2t}\|^{2}]\lesssim d$ and $\mathbb{E}[|\boldsymbol{\theta}^{\top}\boldsymbol{g}^{2t}|\|\boldsymbol{g}^{2t}\|^{2}]\lesssim d$, which implies that the expectation of the last two terms of (B.32) is bounded by $\lesssim\kappa^{2t}c_{\eta}^{2}c_{\xi}^{2}d^{-(p_{*}-1)\lor 1}\lor c_{\eta}^{3}c_{\xi}^{3}d^{-(\frac{3p_{*}}{2}-1)\lor 2}\leq c_{\eta}^{2}c_{\xi}^{2}d^{-(p_{*}-1)\lor 1}\kappa^{2t}$.
Now we bound $\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{g}^{2t}]$ by $\gtrsim c_{\eta}^{I-1}\kappa^{p_{*}-1}$. Let $C_{\sigma}$ be the maximum degree of the activation function with non-zero coefficients of Hermite expansion, which may be infinity when we consider general link functions, and there appear some infinite sums. For these cases we simply assume the sums converge – we discuss the validity of this condition in Section B.1.2. We omit the subscript $2t$ in the following for simplicity. We divide the analysis into the two cases.
(I) If $I=1\Leftrightarrow\mathrm{IE}(\sigma_{*})=\mathrm{GE}(\sigma_{*})=p_{*}$.
For the first term of $\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{g}]$, we have
$\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]$ $=\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\mathbb{E}\bigg{[}\bigg{(}\sum_{j=p_{*}}^{\infty}\alpha_{j}\mathsf{He}_{j}(\boldsymbol{\theta}^{\top}\boldsymbol{x})\bigg{)}\bigg{(}\sum_{j=1}^{C_{\sigma}}j\beta_{j}\mathsf{He}_{j-1}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\bigg{)}\boldsymbol{x}\bigg{]}$ (B.33)
$=\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\sum_{j=p_{*}}^{\infty}\bigg{[}j!\alpha_{j}\beta_{j}\big{(}\boldsymbol{\theta}^{\top}\boldsymbol{\omega}\big{)}^{j-1}\boldsymbol{\theta}+(j+2)!\alpha_{j}\beta_{j+2}\big{(}\boldsymbol{\theta}^{\top}\boldsymbol{\omega}\big{)}^{j}\boldsymbol{\omega}\bigg{]}$ (B.34)
$=\sum_{j=p_{*}}^{C_{\sigma}}j!\alpha_{j}\beta_{j}\big{(}\boldsymbol{\theta}^{\top}\boldsymbol{\omega}\big{)}^{j-1}\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{\theta}$ (B.35)
$=p_{*}!\alpha_{p_{*}}\beta_{p_{*}}\kappa^{p_{*}-1}+O(\kappa^{p_{*}}).$ (B.36)
For the second term of $\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{g}]$, the following decomposition can be made.
$\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]$ (B.37)
$=\sum_{i=1}^{C_{\sigma}-1}(i!)^{-1}\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\mathbb{E}\bigg{[}y\sigma^{(i+1)}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}y\sigma^{(1)}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{)}^{i}\boldsymbol{x}\bigg{]}+\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]$ (B.38)
$=\sum_{i=1}^{C_{\sigma}-1}(i!)^{-1}\eta^{i}\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\mathbb{E}\bigg{[}\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2i}y^{i+1}\sigma^{(i+1)}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\sigma^{(1)}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{)}^{i}\boldsymbol{x}\bigg{]}$ (B.39)
$\quad+p_{*}!\alpha_{p_{*}}\beta_{p_{*}}\kappa^{p_{*}-1}+O(\kappa^{p_{*}}).$ (B.40)
We evaluate each term in the summation. We need to show that although $\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}$ is a function of $\boldsymbol{\theta}^{\top}\boldsymbol{x}$ and $\boldsymbol{\omega}^{\top}\boldsymbol{x}$, it is mostly independent from the two quantities. To verify this, let $\boldsymbol{e}=\frac{\boldsymbol{\theta}-(\boldsymbol{\theta}^{\top}\boldsymbol{\omega})\boldsymbol{\omega}}{\|\boldsymbol{\theta}-(\boldsymbol{\theta}^{\top}\boldsymbol{\omega})\boldsymbol{\omega}\|}$ be the orthogonal component of $\boldsymbol{\theta}$ to $\boldsymbol{\omega}$. Then, we have that
$\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}$ $=\boldsymbol{x}^{\top}(I-\boldsymbol{\omega}\boldsymbol{\omega}^{\top}-\boldsymbol{e}\boldsymbol{e}^{\top})\boldsymbol{x}+(\boldsymbol{e}^{\top}\boldsymbol{x})^{2}$ (B.41)
$=\underbrace{\boldsymbol{x}^{\top}(I-\boldsymbol{\omega}\boldsymbol{\omega}^{\top}-\boldsymbol{e}\boldsymbol{e}^{\top})\boldsymbol{x}}_{\sim\chi_{d-2}^{2},\text{ independent from $\boldsymbol{\omega}^{\top}\boldsymbol{x}$ and $\boldsymbol{\theta}^{\top}\boldsymbol{x}$}}+\bigg{(}\frac{\boldsymbol{\theta}^{\top}\boldsymbol{x}-(\boldsymbol{\theta}^{\top}\boldsymbol{\omega})\boldsymbol{\omega}^{\top}\boldsymbol{x}}{\|\boldsymbol{\theta}-(\boldsymbol{\theta}^{\top}\boldsymbol{\omega})\boldsymbol{\omega}\|}\bigg{)}^{2}.$ (B.42)
With $P_{\boldsymbol{\omega},\boldsymbol{\theta}}=\boldsymbol{I}_{d}-\boldsymbol{\omega}\boldsymbol{\omega}^{\top}-\boldsymbol{e}\boldsymbol{e}^{\top}$, (B.39) is expanded as